OpenStudy (cenaida):

Solve the system, using substitution. What is the n coordinate of the solution? 2m - 3n = 6 and m = 3 - 2n A. 2 B. 0 C. -4 D. 2 E. no solution

OpenStudy (john_es):

\[2(3-2n)-3n=6\Rightarrow 6-4n-3n=6\Rightarrow-7n=6-6\Rightarrow-7n=0\Rightarrow n=0\]

OpenStudy (cenaida):

Thanks again John! I really appreciate you helping me :)

OpenStudy (anonymous):

So we have, two equations. Usually, so include a general solution: with two linear equations in the form of \(y=mx+b\), one can isolate like so: \[\eqalign{ &1.\phantom{space}y=s_1x+b_1 \\ &2.\phantom{space}y=s_2x+b_2 \\ &\\ &s_1x+b_1=s_2x+b_2 \\ &(s_1-s_2)x=b_2-b_1 \\ &\\ &x=\frac{b_2-b_1}{s_1-s_2} \\ }\] So let us rearrange both equations into this form: \[\eqalign{ &2m-3n=6 \\ &2m=6+3n \\ &m=\frac{3n+6}{2} \\ &m=(1.5)n+3 \\ } \phantom{space}\eqalign{ &m=3-2n \\ &m=(-2)n+3 }\] Therefore, we can derive the following values: \[\eqalign{ &s_1=1.5 \\ &b_1=3 \\ &s_2=-2 \\ &b_2= 3 \\ }\] So we can plug into equation: \[n=\frac{b_2-b_1}{s_1-s_2}=\frac{3-3}{1.5-(-2)}=\frac{0}{1.5+2}=\frac{0}{3.5}=0\] Therfore, \(n=0\)